r(m) = 0.36 k(N/m) = 414 m(kg) = 23 x1(m)= 0.350 x2(m)=0.350 x3(m) = 0.700 Q0(deg)=50 Q1(deg)=7 ()0 (rad/s)=2 () (rad/s)=74. (25 points) The lever OĂ, which has mass m and radius of gyration r about O, has an angular velocity w = w, when 0 = 0. Th
Answer #1

Position 1 Position 2 l2 sw. chao Thao Δω. hi at 0 WE, Iw?s my?w3. 13/09 (2) 2 K.E = 5.9616 J P. Es = mgho & Kn хо xo = le-liPEA = (23)9:301 0.35 sin so) + 41460-18047 P.EL= 67. 231 J. at Positiona K. Edga I w? = 022(0.36373? 73.0296 J P. Ez = mgh, +Please give a thumbs up